# Applications of higher order Markov models and Pressure Index to strategize controlled run chases in Twenty20 cricket

### v1

@@ Abstract
In limited overs cricket, the team batting first posts a target score for the team batting second to achieve in order to win the match. The team batting second is constrained by decreasing resources in terms of number of balls left and number of wickets in hand in the process of reaching the target as the second innings progresses. The Pressure Index, a measure created by researchers in the past, serves as a tool for quantifying the level of pressure that a team batting second encounters in limited overs cricket. Through a ball-by-ball analysis of the second innings, it reveals how effectively the team batting second in a limited-over game proceeds towards their target. This research employs higher order Markov chains to examine the strategies employed by successful teams during run chases in Twenty20 matches. By studying the trends in successful run chases spanning over $16$ years and utilizing a significant dataset of $3378$ Twenty20 matches, specific strategies are identified. Consequently, an efficient approach to successful run chases in Twenty20 cricket is formulated, effectively limiting the Pressure Index to $[0.5,3.5]$ or even further down under $0.5$ as early as possible. The innovative methodology adopted in this research offers valuable insights for cricket teams looking to enhance their performance in run chases.
@@ 1 Introduction
Cricket is a team sport played between two teams consisting of 11 players each, following a set of rules as per the format of the game being Tests, One Day Internationals or Twenty20 Internationals (First Class, List A and Twenty20 respectively on a larger basis). In addition, shorter formats of the game like T10, The Hundred, etc., are being introduced for rapid spread and popularity of the game worldwide. One of the major reasons for the intense popularity of cricket nowadays is the uncertainty during a game and it has been observed that the shorter the format, the higher the uncertainty. The above claim is in par with the fact that in the $5$-year period from July $2014$ to June $2019$, $29.89\ \%$ Test matches have been won by teams outside the top $3$ ICC men’s Test rankings against teams ranked within the top $3$ ICC men’s Test rankings, which increases to $33.67\ \%$ if we consider Twenty20 cricket during the same period. The rise is even more in the following $5$-year period from July $2019$ to June $2024$ where it increases from $19.48\ \%$ in Tests to $29.30\ \%$ in Twenty20 cricket.  [^3] Source of data: https://stats.espncricinfo.com/ci/engine/stats/index.html
@@ 2 Review of literature
Though the paper is heavily dependent on the Pressure Index for formulating a strategic approach towards successful run chases in the Twenty20 format of cricket, the work of [Bhattacharjee and Lemmer, 2016] and their subsequent extended works, to name a few, [Talukdar and Bhattacharjee, 2018] and [Bhattacharjee et al., 2018] are not the only contributions in this domain of tactic building for run chases.
@@
Some earlier works that were concerned with the progress of scoring in the second innings of limited overs matches are [Bailey and Clarke, 2006] and [De Silva et al., 2001] While [Bailey and Clarke, 2006] attempted to predict the final score of the second innings of a One-day match while the game was in progress, [De Silva et al., 2001] attempted to estimate the magnitude of victory in One-day matches. In both cases the authors used the Duckworth–Lewis resource table at the end of each over of the second innings of the match.
@@
[Choudhury and Bhargava, 2007] used Multilayer Perceptron Neural Networks to predict the outcome of a cricket tournament involving more than two teams. [Sankaranarayanan, 2014] developed a prediction system that takes in historical ODI cricket match data as well as the instantaneous state of a match and predicts the score at key points in the future, culminating in a prediction of victory or loss. Using factors of the current situation of a cricket match, [Lokhande and Chawan, 2018] attempted to predict the outcome through a two-stage process. In the first stage, the match situation and other cricketing attributes are used to predict the first innings score and then in the second stage the possibility of the team batting second achieving the target is estimated. [Asif and McHale, 2016] presented a model for forecasting the outcomes of One-Day International cricket matches when the game is in progress using dynamic logistic regression. [O’Riley and Ovens, 2006] compared four different score prediction tools by extending them to a ball-by-ball prediction model for predicting the final score, based on the present state of the first innings of a match. A series of papers are available in the literature that used different machine learning tools for predicting either the target of the match or the winner of the match when the match is in progress, viz. [Pathak and Wadhwa, 2016], [Passi and Pandey, 2018], [Vistro et al., 2019], [Modekurti, 2020]. Another set of researchers worked on the development of the Win and Score Predictor (WASP). The purpose of WASP is to predict the score at the end of the first innings of a limited overs match given the current situation and to predict the probability of a successful run chase of the team batting second given the target. Some literature concerning this includes [Hogan et al., 2017], [Singh et al., 2015], and [Nimmagadda et al., 2018].
3 Objectives
@@
    To study the patterns of such run chases and accordingly prepare a strategy in order to keep the team batting second in control throughout a run chase for future Twenty20 matches.
@@ 4.1 Markov chain
Definition 4.1.1  A random process $\{X_{n}:n\geq 0\}$ is said to be a Markov chain with state space $S$ if it satisfies the Markov condition, i.e., for any states $x_{0}$, $x_{1}$, $x_{2}$, …, $x_{n-1}$, $i$, $j\in S$,
@@
Definition 4.1.2  A transition matrix $\mathbf{P}=((p_{ij}))$ is the matrix whose $(i,j)$th entry is given by
@@ 4.2 Markov chains of higher order
Definition 4.2.1  A random process $\{X_{n}:n\geq 0\}$ is said to be a Markov chain of order $k$ if for any states $x_{0}$, $x_{1}$, $x_{2}$, …, $x_{n-1}$, $i$, $j\in S$,
@@
For our purpose, consider $\{X_{n}:n\geq 1\}$ to be the sequence of Pressure Index values (detailed in Section 5) for a team batting second at the end of each over during a run chase in a Twenty20 match, i.e., $X_{n}$ is the Pressure Index value at the end of the $(n+1)$th over during a run chase in a Twenty20 match. We will see why such a sequence is a suitable choice for Markov modelling.
@@ 5 Introduction to Pressure Index
5.1 Pressure Index
As mentioned earlier, the Pressure Index (PI) is a ball-by-ball measure which quantifies the pressure experienced by the team batting second. In Twenty20 cricket, the team batting second has to achieve a certain target before its resources are exhausted. Thus, the pressure of the team batting second reduces if runs are scored at a fast rate and if a considerable number of wickets are kept in hand until the target is reached. Based on this idea, the formula of the Pressure Index [Bhattacharjee and Lemmer, 2016] for teams batting second in Twenty20 matches is given by:
@@ 5.1 Pressure Index
$CRRR$ denotes the current required run rate at any point during the second innings. If $R^{\prime}$ runs have been scored in $B^{\prime}$ balls at some stage, then
@@
The term $\sum w_{i}$ denotes the sum of the weights of the wickets that have fallen until any stage of the innings. As wickets fall, the team’s wicket strength deteriorates. For example, if a top order batter gets out, the wicket strength of the team drops more than when a lower order batter gets out. In order to account for the varying abilities of batters affecting their wicket strengths, wicket weights ($w_{i}$) as per [Lemmer, 2015] are used instead of assigning equal weights to every wicket. The wicket weights for different batting positions, as assigned by [Lemmer, 2015], are given below.
@@
Several other formulae to calculate the Pressure Index are available but we will avoid using those in this study due to various limitations. One such formula, as suggested by [Shah and Shah, 2014], is as follows:
@@
Considering all these factors, the formula given in 6 seems to be the most commendable, and we will consider the same for further study in this paper.
5.2 Working of the Pressure Index
It is evident from 6 that the Pressure Index value of a team batting second remains at unity at the start of a run chase. With every delivery bowled, the value might change and it gradually reaches zero once the team has achieved its target. In case the team fails to do so, the PI value rises steeply at the later stages of the match. To better understand how the indices work, the example of two matches have been taken.
Consider the third Twenty20 match played between South Africa and India at Newlands Cricket Ground, Cape Town on February 24, 2018. South Africa started slowly but cautiously, while chasing 173, being $45/1$ after 9 overs. Although the required run rate was creeping up, 9 wickets were still in hand and they were still in the game. But just when they started to accelerate, wickets began falling, which resulted in a halt in their scoring rate leading to a rapid increase in the Pressure Index. By the 18th over, SA completely lost their grip on the run chase and finally fell short by 7 runs even after scoring 45 runs off their last 3 overs. On the other hand, Australia succesfully chased down a record 244 runs just a week earlier at Auckland with the strategy of never letting the required runrate rise too high.
@@ 5.2 Working of the Pressure Index
Note: The dots in the Pressure curves denote fall of wickets
@@ 6.1 Estimating the order of Markov chain
We have defined our sequence to be the chronological $PI$ values at the end of each over of the second innings of a Twenty20 match but the number of preceding $PI$ values on which the $PI$ at a certain stage of the sequence depends is still unknown. In an ideal scenario, the runs scored and wickets lost in a certain over should technically depend on that in each of the previous overs in the innings. However, in order to avoid unnecessarily complex calculations and to achieve an optimum level of accuracy, one should try to find out the desired order of the Markov chain, i.e. the desired number of preceding $PI$ values on which the next $PI$ value in a sequence depends, that fits our data and leads to a high enough precision for prediction purposes.
@@
All men’s Twenty20 matches in which the team batting second have either reached the target after batting for more than 18 overs or failed to reach the target and have lost the match by a margin not more than 10 runs have been considered as the training sample for the study. Since the goal of our study is to prepare a strategy in support of teams batting second in Twenty20 matches, a certain section of the selected sample of matches will be helpful in replicating the patterns of successful run chases. The choice of such selection has also been made based on the idea of using a significantly large training sample in order to enhance the construction of the model.

### v2

@@ Abstract
In limited overs cricket, the team batting first posts a target score for the team batting second to achieve in order to win the match. The team batting second is constrained by decreasing resources in terms of number of balls left and number of wickets in hand in the process of reaching the target as the second innings progresses. The Pressure Index, a measure created by researchers in the past, serves as a tool for quantifying the level of pressure that a team batting second encounters in limited overs cricket. Through a ball-by-ball analysis of the second innings, it reveals how effectively the team batting second in a limited-over game proceeds towards their target. This research employs higher order Markov chains to examine the strategies employed by successful teams during run chases in Twenty20 matches. By studying the trends in successful run chases spanning over $16$ years and utilizing a significant dataset of $6537$ Twenty20 matches, specific strategies are identified. Consequently, an efficient approach to successful run chases in Twenty20 cricket is formulated, effectively limiting the Pressure Index to $[0.5,3.5]$ or even further down under $0.5$ as early as possible. The innovative methodology adopted in this research offers valuable insights for cricket teams looking to enhance their performance in run chases.
@@ 1 Introduction
Cricket is a team sport played between two teams consisting of 11 players each, following a set of rules as per the format of the game being Tests, One Day Internationals or Twenty20 Internationals (First Class, List A and Twenty20 respectively on a larger basis). In addition, shorter formats of the game like T10, The Hundred, etc., are being introduced for rapid spread and popularity of the game worldwide. One of the major reasons for the intense popularity of cricket nowadays is the uncertainty during a game and it has been observed that the shorter the format, the higher the uncertainty. The above claim is consistent with the fact that in the $5$-year period from July $2014$ to June $2019$, 29.89%29.89\ $\%$ Test matches have been won by teams outside the top $3$ ICC men’s Test rankings against teams ranked within the top $3$ ICC men’s Test rankings, which increases to 33.67%33.67\ $\%$ if we consider Twenty20 cricket during the same period. The rise is even more in the following $5$-year period from July $2019$ to June $2024$ where it increases from 19.48%19.48\ $\%$ in Tests to 29.30%29.30\ $\%$ in Twenty20 cricket [^3] The data for the results of these matches have been collected from https://stats.espncricinfo.com/ci/engine/stats/index.html.
@@
1.1 Basic terminologies
A few basic terms prevalant in Twenty20 cricket are described as follows.
  1. i.
    Overs: In cricket, the batsman stands at one end of the pitch, and the bowler delivers the ball from the other end. Now, after the bowler has delivered six consecutive legal balls, he is said to have completed an over.
  2. ii.
    Powerplay: The first $6$ overs of a Twenty20 match are called powerplay, and a maximum of two fielders can be placed outside the 30-yard circle during this period (see [22]).
  3. iii.
    Death Overs: The last few overs of any innings of a cricket match are said to be Death Overs. We will typically denote the last $4$ overs of a Twenty20 innings to be Death Overs throughout the paper.
  4. iv.
    Middle Overs: The overs following the powerplay till the beginning of the death overs is said to be Middle Overs.
  5. v.
    Openers: Since batting takes place in pairs in cricket, the pair of batters who come in to bat at the start of an innings are termed as Opening partners or simply, openers (see [22]).
  6. vi.
    Batting Depth: If the bowlers of a team, who generally come to bat lower down the order, also possess good batting skills, then the team is said to have significant batting depth.
  7. vii.
    Tailenders: Players who are participating in the team primarily for bowling skills and aren’t a master in the art of batting are called Tailenders or late-order batters (see [8]).
  8. viii.
    Target: In limited overs cricket, the team batting first (refer as Team $A$) tries to score as many runs as they can in the allocated number of overs, without losing all their wickets. At the end of the innings, one run more than the total number of runs scored by Team $A$ is the target for the opposition (Team $B$, say). Team $B$ has to reach the target runs within the alloted number of overs in order to win the match.
Following a literature review and discussions on the novelty of this work in Section 2, in Section 3, we present the fundamental concepts of Markov chains and motivate the need for higher-order models in the context of run-chase dynamics. Section 4.1 introduces the Pressure Index in detail, outlines its computation, and demonstrates how its over-wise evolution reflects match situations. In Section 5.1, we estimate the optimal order of the proposed Markov model. In Section 5.2, we discuss discretization of Pressure Index suggesting suitable gridding. Sections 5.3 to 5.5 describe the construction of transition matrices, distributional modelling of PI across phases of an innings, and contextual factors such as home/away conditions and tournament-specific behaviour. Building on these results, Section 6 develops a prescriptive framework that provides situation-specific Pressure Index targets and strategic guidelines for teams aiming to control run chases in Twenty20 cricket.
@@ 2 Review of literature
Though the paper is heavily dependent on the Pressure Index for formulating a strategic approach towards successful run chases in the Twenty20 format of cricket, the work of [4] and their subsequent extended works, to name a few, [21] and [3] are not the only contributions in this domain of tactic building for run chases.
@@
Some earlier works that were concerned with the progress of scoring in the second innings of limited overs matches are [2] and [6] While [2] attempted to predict the final score of the second innings of a One-day match while the game was in progress, [6] attempted to estimate the magnitude of victory in One-day matches. In both cases the authors used the Duckworth–Lewis resource table at the end of each over of the second innings of the match.
@@
[5] used Multilayer Perceptron Neural Networks to predict the outcome of a cricket tournament involving more than two teams. [18] developed a prediction system that takes in historical ODI cricket match data as well as the instantaneous state of a match and predicts the score at key points in the future, culminating in a prediction of victory or loss. Using factors of the current situation of a cricket match, [11] attempted to predict the outcome through a two-stage process. In the first stage, the match situation and other cricketing attributes are used to predict the first innings score and then in the second stage the possibility of the team batting second achieving the target is estimated. [1] presented a model for forecasting the outcomes of One-Day International cricket matches when the game is in progress using dynamic logistic regression. [14] compared four different score prediction tools by extending them to a ball-by-ball prediction model for predicting the final score, based on the present state of the first innings of a match. A series of papers are available in the literature that used different machine learning tools for predicting either the target of the match or the winner of the match when the match is in progress, viz. [16], [15], [23], [12]. Another set of researchers worked on the development of the Win and Score Predictor (WASP). The purpose of WASP is to predict the score at the end of the first innings of a limited overs match given the current situation and to predict the probability of a successful run chase of the team batting second given the target. Some literature concerning this includes [7], [20], and [13].
Most of the available literature regarding cricket match situation analysis are directed towards predicting final scores, match outcomes, or win probabilities using traditional or recent statistical or data mining tools. The models applied and comprehensive methodologies developed by previous researchers are highly valuable for forecasting. But such models are either descriptive or predictive in nature. However, in this current study, the researchers tried to develop a prescriptive model. By combining a measure of quantifying match pressure with a complex third-order Markov model, the current research attempts to frame strategies for successful run chases in Twenty20 cricket. The approach provides an exclusive, real-time tool for coaches to manage the pressure during run-chase on an over-by-over basis, which differentiates the contribution of the paper from the existing body of literature.
2.1 Objectives
@@
    To study the patterns of run chases and accordingly prepare strategies in order to keep teams batting second in control throughout run chases for future Twenty20 matches.
@@ 3.1 Markov chain
Definition 3.1.
A random process $\{X_{n}:n\geq 0\}$ is said to be a Markov chain with state space $S$ if it satisfies the Markov condition, i.e., for any states $x_{0}$, $x_{1}$, $x_{2}$, …, $x_{n-1}$, $i$, $j\in S$,
@@
Definition 3.2.
A transition matrix $\mathbf{P}=((p_{ij}))$ is the matrix whose $(i,j)$th entry is given by
@@ 3.2 Markov chains of higher order
Definition 3.3.
A random process $\{X_{n}:n\geq 0\}$ is said to be a Markov chain of order $k$ if for any states $x_{0}$, $x_{1}$, $x_{2}$, …, $x_{n-1}$, $i$, $j\in S$,
@@
For our purpose, consider $\{X_{n}:n\geq 1\}$ to be the sequence of Pressure Index values (see details in Section 4) for a team batting second at the end of each over during a run chase in a Twenty20 match, i.e., $X_{n}$ is the Pressure Index value at the end of the $(n+1)$th over during a run chase in a Twenty20 match. We will see why such a sequence is a suitable choice for Markov modelling.
@@ 4 Introduction to Pressure Index
4.1 Definitions and Formulation
As mentioned earlier, the Pressure Index ($PI$) is a ball-by-ball measure which quantifies the pressure experienced by the team batting second. In Twenty20 cricket, the team batting second has to achieve a certain target before its resources are exhausted. Thus, the pressure of the team batting second reduces if runs are scored at a fast rate and if a considerable number of wickets are kept in hand until the target is reached. Based on this idea, the formula of the Pressure Index [4] for teams batting second in Twenty20 matches is given by:
@@ 4.1 Definitions and Formulation
CRRRCRRR denotes the current required run rate at any point during the second innings. If $R^{\prime}$ runs have been scored in $B^{\prime}$ balls at some stage, then
@@
The term $\sum w_{i}$ denotes the sum of the weights of the wickets that have fallen until any stage of the innings. As wickets fall, the team’s wicket strength deteriorates. For example, if a top order batter gets out, the wicket strength of the team drops more than when a lower order batter gets out. In order to account for the varying abilities of batters affecting their wicket strengths, wicket weights ($w_{i}$) as per [9] are used instead of assigning equal weights to every wicket. The wicket weights for different batting positions, as assigned by [9], are given below.
Table 1: Wicket weights assigned by Lemmer
@@
Several other formulae to calculate the Pressure Index are available but we will avoid using those in this study due to various limitations. One such formula, as suggested by [19], is as follows:
@@
Considering all these factors, the formula (6) seems to be the most commendable, and we will consider the same for further study in this paper.
4.2 Operational Mechanism
It is evident from (6) that the Pressure Index value of a team batting second remains at unity at the start of a run chase. With every delivery bowled, the value might change and it gradually reaches zero once the team has achieved its target. In case the team fails to do so, the PI value rises steeply at the later stages of the match. To better understand how the indices work, the example of two matches have been taken.
Consider the Twenty20 match played between South Africa and India at Newlands Cricket Ground, Cape Town on February 24, 2018. South Africa started slowly but cautiously, chasing 173, $45/1$ after 9 overs. Although the required run rate was creeping up, 9 wickets were still in hand and they were still in the game. But just when they started to accelerate, wickets began falling, which resulted in a halt in their scoring rate leading to a rapid increase in the Pressure Index. By the 18th over, SA completely lost their grip on the run chase and finally fell short by 7 runs even after scoring 45 runs off their last 3 overs. On the other hand, Australia successfully chased down a record 244 runs, just a week prior to that, at Auckland with the strategy of never letting the required run rate rise too high.
@@ 4.2 Operational Mechanism
(Note: The dots in the pressure curves denote fall of wickets)
@@ 5.1 Estimating the order of Markov chain
We have defined our sequence to be the chronological $PI$ values at the end of each over of the second innings of a Twenty20 match but the number of preceding $PI$ values on which the $PI$ at a certain stage of the sequence depends is still unknown. In an ideal scenario, the runs scored and wickets lost in a certain over should technically depend on that in each of the previous overs in the innings. However, in order to avoid unnecessarily complex calculations and to achieve an optimum level of accuracy, one should try to find out the desired order of the Markov chain, i.e. the desired number of preceding $PI$ values on which the next $PI$ value in a sequence truly depends, that fits our data well and leads to a high enough precision for prediction purposes.
@@
All men’s Twenty20 matches in which the team batting second have either reached the target after batting for more than 18 overs or failed to reach the target and have lost the match by a margin not more than 10 runs have been considered for the study. Since the goal of our study is to prepare a strategy in support of teams batting second in Twenty20 matches, a certain section of the selected sample of matches will be helpful in replicating the patterns of successful run chases. The choice of such selection has also been made based on the idea of using a significantly large training sample in order to enhance the construction of the model.